Starting with a stock solution of 4 mM KMnO4, you plan to make 16 mL of each of the following concentrations of KMnO4: 1.0 mM, 250 μM, 62.5 μM , and 15.6 μM. What is the V1, expressed in mL, for this dilution series? Report your answer to 2 decimal places.
Answer #1

here M1= 4mM and V2=16 mL

1) 1.0 mM

M1V1=M2V2

V1= M2V2/M1

= 1*16/4

=4 mL

2)250 μM

1 uM= 10^-3 mM

250 *10^-3 mM

M1V1=M2V2

V1= M2V2/M1

= 250*10^-3*16/4=1

V1= 1mL

3) 62.5 μM

1 uM= 10^-3 mM

62.5*10^-3 mM

M1V1=M2V2

V1= M2V2/M1

= 62.5*10^-3*16/4=1/4=0.25

V1= 0.25 mL=250uL ( 1 mL=1000 uL)

4) 15.6 μM

1 uM= 10^-3 mM

15.6*10^-3 mM

M1V1=M2V2

V1= M2V2/M1

= 15.6*10^-3*16/4=0.0624mL

V1=0.0624 *1000=62.4 uL